VDSL Company has two mutually exclusive projects. Below is a table representing the initial investment and...

  1. VDSL Company has two mutually exclusive projects. Below is a table representing the initial investment and cash flows for these projects over four (4) years.

Project A

Project B

Year

Cash Flow

Cash Flow

$

$

0

-750,000

-750,000

1

250,000

200,000

2

350,000

400,000

3

250,000

100,000

4

200,000

175,000

a. If the company’s required rate of return is 8%, calculate the Profitability Index of each project and determine which project is the best investment.

b.   If the company wishes to recover the investment in 2.5 years, calculate the payback period of each project and determine which project is the best investment.                    
Answer #1
a. Project A Project B
Profitability Index 1.17 0.98
Project A is the best investment as it has profitability index of more than 1.
Working:
# 1: Calculation of present value of cash inflow
Project A Project B
Year Discount factor Cash inflow Discounted cash inflow Cash inflow Discounted cash inflow
a b=1.08^-a c d=c*b e f=e*b
1 0.925926 $   2,50,000.00 $   2,31,481.48 $    2,00,000.00 $   1,85,185.19
2 0.857339 $   3,50,000.00 $   3,00,068.59 $    4,00,000.00 $   3,42,935.53
3 0.793832 $   2,50,000.00 $   1,98,458.06 $    1,00,000.00 $       79,383.22
4 0.73503 $   2,00,000.00 $   1,47,005.97 $    1,75,000.00 $   1,28,630.22
Total $   8,77,014.10 $   7,36,134.16
# 2:Calculation of profitability index
Profitability index = Present value of cash inflow / Investment
So,
Profitability index of:
Project A = $   8,77,014.10 / $ 7,50,000.00 =                      1.17
Project B = $   7,36,134.16 / $ 7,50,000.00 =                      0.98
b. Project A Project B
Payback Period 2.60 Years 3.29 Years
Although both project failed to recover the investment in 2.5 years, Project A has earlier recovery than Project B.
So, Project A is best investment.
Working:
Project A Project B
Year Cash flow Cumulative cash flow Cash flow Cumulative cash flow
0 $ -7,50,000.00 $ -7,50,000.00 $ -7,50,000.00 $ -7,50,000.00
1 $   2,50,000.00 $ -5,00,000.00 $    2,00,000.00 $ -5,50,000.00
2 $   3,50,000.00 $ -1,50,000.00 $    4,00,000.00 $ -1,50,000.00
3 $   2,50,000.00 $   1,00,000.00 $    1,00,000.00 $     -50,000.00
4 $   2,00,000.00 $   3,00,000.00 $    1,75,000.00 $   1,25,000.00
Payaback period of:
Project A = 2+(150000/250000) =                    2.60
Project B = 3+(50000/175000) =                    3.29